Assume that yo-yo drops 1.0 m vertically from its owner's hand at a uniform speed of 0.80 m/s along its string. At the end of the string, it changes direction in 0.10 s and rises toward the owner's hand at a uniform speed of 0.60 m/s. Compute the average acceleration of the yo-yo at the bottom end of the string as it changes direction.


1
Expert's answer
2022-03-29T12:29:20-0400

Answer

acceleration is that

"a=\frac{v-v'}{t-t'}\\=\frac{0.8+0.6}{0.1}\\=14m/s^2"